### v1

# Ground-Compose-Reinforce: Tasking Reinforcement Learning Agents through Formal Language

Abstract
Grounding language in complex perception (e.g. pixels) and action is a key challenge when building situated agents that can interact with humans via language. In past works, this is often solved via manual design of the language grounding or by curating massive datasets relating language to elements of the environment. We propose Ground-Compose-Reinforce, a neurosymbolic framework for grounding formal language from data, and eliciting behaviours by directly tasking RL agents through this language. By virtue of data-driven learning, our framework avoids the manual design of domain-specific elements like reward functions or symbol detectors. By virtue of compositional formal language semantics, our framework achieves data-efficient grounding and generalization to arbitrary language compositions. Experiments on an image-based gridworld and a MuJoCo robotics domain show that our approach reliably maps formal language instructions to behaviours with limited data while end-to-end, data-driven approaches fail.
1 Introduction
Grounded language acquisition—the way a situated agent associates language with perception and action in its environment—is a fundamental problem when building robots and other AI systems that are interfaced through language. In recent years, this problem has garnered significant interest from the AI community. One popular solution is to start with manually-designed groundings of language based on domain knowledge, such as a language-conditional reward function or success detector (e.g. [1, 2, 3, 4]). For instance, in the BabyAI benchmark [1], successful execution of instructions like “go to the red ball” can be evaluated programmatically via the environment simulator, providing a reward signal for learning language-conditioned behaviours. Such language groundings generalize to arbitrary scenarios and controlled subsets of language by design, but are hard to hand-engineer in complex, non-simulated settings based on raw perceptual inputs like pixels.
The recent advent of large language models has inspired an alternative, scalable solution to grounded language acquisition: training on diverse datasets pairing language descriptions with environment trajectories (e.g. $\pi_{0}$ [5], RT-2 [6], LIV [7], VPT [8]). While this requires no manual design of the language grounding, enormous datasets are needed to capture the intricacies of language and its corresponding range of behaviours [1, 9, 10]. Even then, fully data-driven language-grounding models are known to struggle on long-horizon or out-of-distribution instructions [11, 12, 13].
To address these issues, we propose a neurosymbolic framework called Ground-Compose-Reinforce for grounded language learning and eliciting diverse behaviours through this language. We hypothesize that compositionality plays a critical role in data-efficient and generalizable grounded language learning, by requiring only the grounding of basic concepts which can then be systematically combined to interpret complex language. To explore this hypothesis, we consider formal languages, whose syntaxes and semantics are defined by compositional rules based on an initial set of atomic symbols. Starting with no prior knowledge except for these rules, our framework: (1) grounds these atomic symbols in the environment through data; (2) composes these symbols via a formal language to express complex tasks; (3) trains an agent to solve any such task in the environment via reinforcement learning (RL) on self-generated rewards based on its grounded language understanding.
We implement this framework for Reward Machines [14], which express reward functions via temporal and logical composition over a set of binary environment properties. Given a small grounding dataset of trajectories labelled with these properties, we show how an agent can learn to ground any Reward Machine in the environment, allowing it to self-evaluate task success. To address sparse rewards, we propose a compositional reward shaping strategy that allows the agent to directly predict optimal value functions for any such task. We evaluate our agent across diverse Reward Machine tasks, including temporally-extended navigation in a 2-D gridworld and robotic manipulation in MuJoCo with raw perceptual inputs and low-level actions. We show that our approach elicits a diversity of behaviours expressible as Reward Machines, outperforming non-compositional grounding approaches according to a ground-truth interpretation of each task. Critically, we show that our approach generalizes to out-of-distribution behaviours that never appear in the grounding dataset and scales to sparse-reward settings, all while requiring no domain-specific engineering of the language grounding (e.g. of success detectors, reward functions, or symbol evaluators). More broadly, our findings highlight the important role of compositionality in grounded language learning.
2 Related Work
Grounded Language Learning. Several past works have explored grounded language learning. Hermann et al. [2] show that an RL agent in a 3-D environment can consistently navigate to target objects described via language by being rewarded for successful trajectories. Hill et al. [3] show that an agent can learn new word-object bindings and apply that knowledge to solve tasks within the same episode. Liu et al. [15] show that a meta-RL agent indirectly learns to interpret language-based advice embedded within an environment to improve its task performance. Chaplot et al. [4] propose a neural architecture for grounded language learning in a 3-D Doom environment, considering both RL with success-based rewards and imitation learning from an oracle policy. Chevalier-Boisvert et al. [1] propose the BabyAI platform for learning a synthetic language in a 2-D gridworld, finding that existing RL and imitation learning approaches are sample inefficient and generalize poorly. Unlike our approach, these methods require significant manual design of reward functions or oracle policies.
Recent works have proposed to learn the language grounding from diverse data, without relying on domain-specific manual design. Black et al. [5], Brohan et al. [6] and Kim et al. [16] propose vision-language-action models for real-world robotics, while Baker et al. [8] and Lifshitz et al. [13] propose similar models for playing MineCraft. Bahdanau et al. [17] and Ma et al. [7] propose approaches for learning language-conditioned reward functions directly from data. Works have also proposed directly leveraging vision-language models for rewards in vision-based environments [18, 19, 20]. These approaches do not explicitly leverage compositional language structure, and are known to struggle or hallucinate on complex, long-horizon tasks. Shi et al. [12], Yuan et al. [21], Huang et al. [22] and Ahn et al. [23] propose to use language models to decompose complex tasks at execution time. This shares motivation with our work, except we instead leverage formal language semantics to decompose tasks.
Formal Languages for RL. Formal languages like Linear Temporal Logic (LTL) [24] and Reward Machines (RMs) [14] have a rich history of application towards the control [25, 26], verification [27, 28, 29], and specification of dynamical systems. More recently, they have risen in popularity in deep RL for white-box specification of reward functions [30, 31, 32, 33] and in many cases, natural language commands can be automatically translated into such formal languages [34, 35, 36, 37]. Several works show that formal languages enable compositional generalization to unseen instructions. Vaezipoor et al. [38], Kuo et al. [39] and Yalcinkaya et al. [40] train instruction-conditioned policies that zero-shot generalize to unseen instructions by training on procedurally generated LTL formulas. Qiu et al. [41], Liu et al. [42], León et al. [43] and Jackermeier and Abate [44] train transferable skills that can be invoked by a planner. Nangue Tasse et al. [45] consider how value functions and policies can be composed zero-shot for arbitrary RM tasks, but assume that all the tasks can be captured via a compact set of “goal states” that is known and enumerable. Our compositional reward shaping approach also builds on prior methods. Camacho et al. [46], Furelos-Blanco et al. [47] and Parać et al. [48] provide potential-based rewards for RMs based on the current RM state, but such methods provide no signal when the propositions themselves rarely occur. Several works propose continuous progress signals for each proposition [49, 50, 51, 52, 53, 54], but these approaches are limited to tasks with a binary success criteria. While progress signals are typically manually specified, we show how they can be learned directly from data.
Nearly all formal-language-based deep RL methods are implemented with manually designed symbol evaluators (a.k.a. a labelling function) based on the environment simulator. To our knowledge, only a few works specifically avoid this assumption, but they instead depend on a manually designed reward signal. Li et al. [55, 56] consider the implications of noisy symbol groundings trained from data when using RMs. Hyde and Santos [57] and Christoffersen et al. [58] infer RMs from the reward signal as an inductive bias for RL. Umili et al. [59, 60], Kuo et al. [39], Andreas et al. [61] and Oh et al. [62] show that ungrounded formal language specifications can improve RL by providing information about the task structure. In contrast to these works, our approach does not rely on manually designed symbol evaluators or reward functions.
3 Preliminaries
3.1 Reinforcement Learning
A reinforcement learning (RL) problem considers an environment modelled as a Markov Decision Process (MDP) $\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{P},R,\mu,\gamma\rangle$, where $\mathcal{S}$ is a set of states, $\mathcal{A}$ a set of actions, $\mathcal{T}\subset\mathcal{S}$ a set of terminal states, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\to\Delta\mathcal{S}$ a transition probability distribution, $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbbm{R}$ a reward function, $\mu\in\Delta\mathcal{S}$ an initial probability distribution, and $\gamma\in[0,1]$ a discount factor. An episode begins with $s_{0}\sim\mu$, and at each time $t\geq 0$ the agent chooses an action $a_{t}$, then observes the next state $s_{t+1}\sim\mathcal{P}(s_{t},a_{t})$ and reward $r_{t+1}=R(s_{t},a_{t},s_{t+1})$, repeating until a terminal state $s_{T}\in\mathcal{T}$ is reached. We refer to full episodes as trajectories, denoted by $\tau=\langle s_{0},a_{0},s_{1},a_{1},\ldots\rangle$. A history $h_{t}=\langle s_{0},a_{0},s_{1},a_{1},\ldots s_{t}\rangle$ refers to the states and actions up to time $t$. The agent’s goal is to interact with the environment to learn a policy $\pi(a_{t}|s_{t})$ that maximizes the expected discounted return $\mathbb{E}_{\tau\sim\pi}[\sum_{t=1}^{T}{\gamma^{t}r_{t}}]$ (where the episode length $T$ can be $\infty$).
Figure 1: Four temporally extended tasks in a simple gridworld expressed as Reward Machines over the propositional vocabulary $\mathcal{AP}=\{\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}},\mathrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}G}},\mathrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}},\triangle,\bigcirc\}$. Edges are labelled with $\langle\varphi,r\rangle$, indicating the logical condition $\varphi$ for when the corresponding transition should be followed, and the reward $r$ that is yielded as a result. Doubled circles indicate terminal states, and we omit non-rewarding self-loop edges to aid readability.
3.2 Task Specification via Reward Machines
Formal languages with sequential or temporal structure—including Reward Machines (RMs) [14], regular expressions [63], and various temporal logics [54, 38]—offer an intuitive and expressive interface for specifying tasks in RL, and support representations that capture high-level task structure and compositionality. Starting with a predefined set of atomic propositions $\mathcal{AP}$ that represent abstract, binary features of environment states, such languages can be used to express temporally extended properties. In this work, we focus on tasks specified as RMs, which subsume several formal languages of interest [46] such as LTL over finite traces [64, 65] and regular expressions.
A Reward Machine (RM) [14] is an abstract automaton that defines a reward function over the set of propositions $\mathcal{AP}$ (Figure 1). It takes as input a sequence of truth assignments $\langle\omega_{1},\omega_{2},\ldots\rangle$ over $\mathcal{AP}$, where each $\omega_{t}\in 2^{\mathcal{AP}}$ denotes the subset of $\mathcal{AP}$ that holds true at time $t$, and outputs a corresponding sequence of rewards $\langle r_{1},r_{2},\ldots\rangle$. An RM has a finite set of internal states $\mathcal{U}$ and begins in a fixed state $u_{0}$ at time $t=0$. At each step $t\geq 1$, the RM updates its state to $u_{t}\in\mathcal{U}$ and emits a reward $r_{t}$ based on the current input $\omega_{t}$ and the previous state $u_{t-1}$. This continues until the RM enters a terminal state from a designated subset $\mathcal{F}\subset\mathcal{U}$.
Running Example.
The gridworld in Figure 1 will serve as a running example. Consider the top-left RM, which describes a task composed of three subgoals to be completed in a fixed order. The input sequence $\langle\omega_{1},\omega_{2},\ldots\rangle$ identifies the values of all propositions at each time $t$: whether the agent is at a red object, a green object, a circle, and so on. The RM state $u_{t}\in\mathcal{U}$ provides memory of which subgoals have been completed, and transitions to the next state as soon as the agent achieves the current subgoal. When a red triangle, a green object, and a blue object that isn’t a triangle are reached in that order, the RM terminates with a reward of 1.
Definition 1.
A Reward Machine (RM) is defined as a tuple $\langle\mathcal{U},u_{0},\mathcal{F},\mathcal{AP},\delta_{u},\delta_{r}\rangle$, where $\mathcal{U}$ is a finite set of states, with initial state $u_{0}\in\mathcal{U}$ and terminal states $\mathcal{F}\subset\mathcal{U}$; $\mathcal{AP}$ is a set of propositions; $\delta_{u}:\mathcal{U}\times 2^{\mathcal{AP}}\to\mathcal{U}$ is a transition function that updates the RM state based on the current truth assignment; and $\delta_{r}:\mathcal{U}\times 2^{\mathcal{AP}}\to\mathbb{R}$ is a reward function that emits a reward at each step.
As shown in Figure 1, the transition and reward functions of an RM can be intuitively and compactly specified via a set of labelled edges of the form $\langle u,u^{\prime},\varphi,r\rangle$, indicating that if a truth assignment $\omega$ satisfies $\varphi$ (denoted by $\omega\models\varphi$), then $\delta_{u}(u,\omega)=u^{\prime}$ and $\delta_{r}(u,\omega)=r$.
3.3 Grounded Interpretations
RMs specify tasks in terms of abstract symbols (e.g., $\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}}$, $\triangle$), but these symbols must be grounded in the environment to be meaningful. This is achieved via a labelling function $\mathcal{L}:\mathcal{S}\to 2^{\mathcal{AP}}$ that maps each MDP state $s$ to the set $\omega\subseteq\mathcal{AP}$ of propositions that hold true in $s$. Any such $\mathcal{L}$ grounds an RM $\mathcal{R}$ over propositions $\mathcal{AP}$, as it induces a reward sequence for any MDP trajectory $\tau$: first, states $s_{t}$ in $\tau$ are converted into truth assignments $\omega_{t}=\mathcal{L}(s_{t})$, then the RM is recursively simulated over the sequence $\langle\omega_{1},\omega_{2},\ldots\rangle$ until termination.
An RM-MDP augments a reward-free MDP with an RM and labelling function. The resulting reward function is generally non-Markovian with respect to the original state space $\mathcal{S}$, since the reward at time $t$ depends on the internal RM state $u_{t}$. One might therefore consider training a history-based policy $\pi(a_{t}|h_{t})$ to act optimally. Fortunately, this is unnecessary if the agent has oracle access to $\mathcal{L}$. In that case, the RM-MDP is Markovian over the extended state space $\mathcal{S}\times\mathcal{U}$ [14]. The agent can use $\mathcal{L}$ to compute $\omega_{t}=\mathcal{L}(s_{t})$ and recursively simulate the RM to track $u_{t}$. This allows training a policy of the form $\pi(a_{t}|s_{t},u_{t})$, where the RM state $u_{t}$ compactly encodes the history $h_{t}$ and is sufficient for optimal decision making.
Definition 2.
An RM-MDP is a triple $\langle\mathcal{M},\mathcal{R},\mathcal{L}\rangle$, where $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mu,\gamma\rangle$ is an MDP without rewards or terminal states, $\mathcal{R}=\langle\mathcal{U},u_{0},\mathcal{F},\mathcal{AP},\delta_{u},\delta_{r}\rangle$ is an RM, and $\mathcal{L}:\mathcal{S}\to 2^{\mathcal{AP}}$ is a labelling function. The RM-MDP is equivalent to an MDP with state space $\mathcal{S}\times\mathcal{U}$ and reward function induced by $\mathcal{R}$ and $\mathcal{L}$.
4 Problem Setting
We aim to develop RL agents that are fully taskable through formal language: given only a formal language specification, it must learn to solve the corresponding task by interacting with the environment. We avoid manually designing reward functions or labelling functions that capture task-relevant aspects of the environment—while these methods conveniently help ground language in the environment, they become increasingly hard to specify as tasks and environments grow in complexity. However, without a way to ground language, any language-based specification is meaningless to an agent acting in an environment. Imagine, for example, that your boss tells you to “get the dax, then throw away the blick[66, 67] but you don’t know what “dax” and “blick” mean.
We explore how formal language can be grounded by training on a fixed grounding dataset, and then subsequently applied for specifying RL tasks. We formulate this problem in the context of RMs.
Problem Formulation. Let $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mu,\gamma\rangle$ be an MDP without rewards or terminal states, and let $\mathcal{AP}$ be a finite set of propositions. We assume a ground-truth labelling function $\mathcal{L}^{*}:\mathcal{S}\to 2^{\mathcal{AP}}$ exists, but is unobservable. Instead, we assume we have a fixed grounding dataset $\mathcal{D}=\{\langle\tau^{i},\omega^{i}\rangle\}_{i=1}^{N}$ of trajectories $\tau^{i}=\langle s^{i}_{0},a^{i}_{0},s^{i}_{1},\ldots\rangle$ with corresponding labels $\omega^{i}=\langle\mathcal{L}^{*}(s^{i}_{0}),\mathcal{L}^{*}(s^{i}_{1}),\ldots\rangle$. We aim to use $\mathcal{D}$ to develop an agent that can learn to solve any RM task $\mathcal{R}=\langle\mathcal{U},u_{0},\mathcal{F},\mathcal{AP},\delta_{u},\delta_{r}\rangle$ over $\mathcal{AP}$ without supervision. That is, it can learn a policy $\pi_{\mathcal{R}}(a_{t}|h_{t})$ that performs well in the corresponding RM-MDP $\langle\mathcal{M},\mathcal{R},\mathcal{L}^{*}\rangle$ by interacting with $\mathcal{M}$, but without access to RM-MDP rewards or $\mathcal{L}^{*}$. [^1] For the problem statement, we consider the more general form of history-based policies, rather than policies conditioned on the ground-truth RM state, which depends on $\mathcal{L}^{*}$.
Running Example.
Suppose our gridworld agent is expected to solve any new RM task over the vocabulary $\{\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}},\mathrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}G}},\mathrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}},\triangle,\bigcirc\}$. The agent may interact freely with the environment, but observes neither ground-truth rewards nor propositional labels (e.g. whether $\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}}$ holds in the current state) while doing so. Instead, it is first given a fixed set $\mathcal{D}$ of trajectories labelled with this vocabulary—from this, it must learn how to interpret any such RM task in the context of the environment.
Since the agent faces arbitrary RM tasks without reward supervision, it must acquire a grounded understanding of both individual symbols (e.g., ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}$) and their compositions as RM tasks to reliably solve new tasks. We also consider sparse settings in which $\mathcal{L}^{*}(s)$ is constant over most of the reachable state space under random exploration. This is common in domains like robotics, where random exploration rarely produces meaningfully diverse or goal-relevant behaviour. We hypothesize that existing RM-based approaches fail to learn effectively under such conditions.
5 Ground-Compose-Reinforce
We address the problem through a framework called Ground-Compose-Reinforce. At a high level, the agent first grounds the symbols of a formal language by training on $\mathcal{D}$. Given any task expressed in this language, it can then evaluate trajectories by leveraging both its grounded interpretation of symbols and the compositional semantics of the formal language (i.e., how symbol meanings combine to form complex tasks). Finally, the agent learns to solve new tasks through RL without supervision, instead using its ability to self-evaluate behaviours to generate its own learning signals.
We hypothesize that grounding language bottom-up—first learning individual symbols, then composing them to interpret complex tasks—is more effective than learning to associate entire task specifications with behaviours in an end-to-end fashion. We test this hypothesis in our experiments.
We now describe the core implementation of Ground-Compose-Reinforce for RMs. This version supports fully unsupervised RL for new RM tasks. In Section 6, we extend this framework to handle sparse settings by learning a compact set of value functions from $\mathcal{D}$ and using them to approximate value functions for any RM task. We show that this enables the agent to solve long-horizon tasks without sacrificing optimality on the target task.
Algorithm 1 Ground-Compose-Reinforce for RMs
0:  MDP $\mathcal{M}$ without rewards, Propositional symbols $\mathcal{AP}$, Dataset $\mathcal{D}$ of labelled trajectories, RM task $\mathcal{R}$ over $\mathcal{AP}$ // Phase 1: Ground symbols via offline learning
1:  Train labelling function ^$\hat{\mathcal{L}}(s)$ on $\mathcal{D}$ using any binary classifier // Phase 2: Solve any RM task using self-generated rewards
2:  Initialize policy $\pi_{\mathcal{R}}(a\mid s,u)$ arbitrarily
3:  for each episode do
4:     Observe initial state $s$ in $\mathcal{M}$; set $u$ to the initial state of $\mathcal{R}$
5:     while $u$ is non-terminal do
6:        Sample action $a\sim\pi_{\mathcal{R}}(\cdot\mid s,u)$
7:        Execute $a$ in $\mathcal{M}$ and observe next state $s^{\prime}$
8:        Compute truth assignment $\hat{\omega}\leftarrow\hat{\mathcal{L}}(s^{\prime})$
9:        Update RM: $u^{\prime}\leftarrow\delta_{u}(u,\hat{\omega})$, $r\leftarrow\delta_{r}(u,\hat{\omega})$
10:        Update $\pi_{\mathcal{R}}$ with RL for transition $\langle s,u,a,r,s^{\prime},u^{\prime}\rangle$
11:        Set $s\leftarrow s^{\prime}$, $u\leftarrow u^{\prime}$
12:     end while
13:  end for
Core Algorithm. We first ground the propositional symbols $\mathcal{AP}$ by training a model of the labelling function ^$\hat{\mathcal{L}}(s)\approx\mathcal{L}^{*}(s)$ on $\mathcal{D}$ using any binary classification method. Then, given any RM task $\mathcal{R}$ over $\mathcal{AP}$, we solve a surrogate RM-MDP $\langle\mathcal{M},\mathcal{R},\hat{\mathcal{L}}\rangle$ via RL. This surrogate task approximates the true RM-MDP $\langle\mathcal{M},\mathcal{R},\mathcal{L}^{*}\rangle$, for which the agent lacks supervision. Although the agent cannot access $\mathcal{L}^{*}$ or ground-truth rewards, it can query $\hat{\mathcal{L}}$ freely—thus, it can simulate rewards $\langle{r}_{1},{r}_{2},\ldots\rangle$ and RM states $\langle{u}_{1},{u}_{2},\ldots\rangle$ for any trajectory, as described in Section 3.3. We use these self-generated signals to train a policy $\pi_{\mathcal{R}}(a_{t}|s_{t},u_{t})$ for the equivalent MDP over state space $\mathcal{S}\times\mathcal{U}$. We outline this full process in Algorithm 1.
Running Example.
With Ground-Compose-Reinforce, the gridworld agent first learns to interpret propositions $\{\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}},\mathrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}G}},\mathrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}},\triangle,\bigcirc\}$ in relation to its environment, grounding each proposition to states via the dataset $\mathcal{D}$. Once grounded, the agent can understand any RM task (e.g. those in Figure 1) that composes these symbols through temporal or logical structure. A human can then specify a new RL task as an RM $\mathcal{R}$, without needing to program a task-specific reward function. The agent learns by predicting its own rewards and RM transitions based on its interpretation of the task.
6 Compositional Reward Shaping
Figure 2: An illustration of how we estimate optimal values in an RM-MDP. Suppose the agent is currently in RM state $u^{A}$ (green and bolded). To evaluate the expected return for the transition $u^{A}\to u^{B}$, we estimate how close the agent is to satisfying the formula on the transition (reaching the red triangle), and bootstrap with a coarse value estimate for RM state $u^{B}$. The overall value of $u^{A}$ is approximated by the maximum expected return across all outgoing transitions from $u^{A}$.
We now address the challenge of sparsity by extending Ground-Compose-Reinforce with potential-based reward shaping [68], using approximations of optimal value functions (OVFs) as shaping potentials. Specifically, we estimate the OVF for the surrogate RM-MDP $\langle\mathcal{M},\mathcal{R},\hat{\mathcal{L}}\rangle$ by decomposing $\mathcal{R}$ into simpler subtasks. The most basic subtasks correspond to satisfying individual propositions or their negations, and we estimate the OVFs for these $2|\mathcal{AP}|$ tasks directly from $\mathcal{D}$ using offline RL. These $2|\mathcal{AP}|$ OVFs then serve as building blocks that can be composed to approximate the OVF for any of the infinitely many RM tasks over $\mathcal{AP}$. We describe this compositional shaping method below.
6.1 Deconstructing RMs into Logical Subtasks
To estimate the OVF of the surrogate RM-MDP $\langle\mathcal{M},\mathcal{R},\hat{\mathcal{L}}\rangle$, denoted $V^{*}_{\mathcal{R}}(s,u)$, we combine two key ideas: (1) treating each RM transition in $\mathcal{R}$ as an independent subtask, and (2) bootstrapping from a state-independent value function $v^{*}_{\mathcal{R}}(u)$.
Consider the example in Figure 2, where the agent is in MDP state $s$ and RM state $u^{A}$. Each RM transition is associated with a propositional condition $\varphi$ (e.g., ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}\land\triangle$ or ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}\land\lnot\triangle$), and we treat the satisfaction of $\varphi$ as its own subtask. Definition 3 formalizes this: we introduce the class $\Diamond\mathbf{PL}(\mathcal{AP})$ of reachability tasks over propositional formulas, where each task $\Diamond\varphi$ entails reaching a state $s_{T}$ such that ^$\hat{\mathcal{L}}(s_{T})\models\varphi$, terminating with reward 1 upon satisfaction. $\Diamond\varphi$ is itself an RM-MDP with a single non-terminal RM state, and is therefore Markovian over $\mathcal{S}$. We denote its OVF as $V^{*}_{\Diamond\varphi}(s)$.
Definition 3.
For any propositional logic formula $\varphi$ over $\mathcal{AP}$, define $\Diamond\varphi$ as the task of reaching a state $s_{T}\in\mathcal{S}$ such that ^$\hat{\mathcal{L}}(s_{T})\models\varphi$. The episode terminates and yields a reward of 1 if such a state is reached, and continues indefinitely with 0 reward otherwise. Let $\Diamond\mathbf{PL}(\mathcal{AP})=\allowbreak\left\{\Diamond\varphi\;\middle|\;\varphi\text{ is a propositional formula over }\mathcal{AP}\right\}$ denote the set of such reachability tasks.
The second component of our method is to bootstrap using a state-independent value function $v^{*}_{\mathcal{R}}(u)$ that approximates the expected return from any RM state $u$ while ignoring the MDP state. We estimate $v^{*}_{\mathcal{R}}(u)$ using a variant of Value Iteration over the RM graph, following Camacho et al. [46]. Finally, to estimate $V^{*}_{\mathcal{R}}(s,u)$, we evaluate each outgoing transition $\langle u,u^{\prime},\varphi,r\rangle$ from $u$ by combining the subtask value $V^{*}_{\Diamond\varphi}(s)$ with the bootstrapped value of the next RM state $v^{*}_{\mathcal{R}}(u^{\prime})$. The final estimate is the maximum expected return over all such transitions:
$V^{*}_{\mathcal{R}}(s,u)\approx\max_{\langle u,u^{\prime},\varphi,r\rangle}\bigl{[}V^{*}_{\Diamond\varphi}(s)\cdot\bigl{(}r+\gamma v^{*}_{\mathcal{R}}(u^{\prime})\bigr{)}\bigr{]}$ (1)
This approximation assumes no RM self-transitions with non-zero rewards. Appendix A provides further justification for our approximation and also extends it to handle any RM.
6.2 Further Deconstructing Logical Subtasks
Approximation 1 allows us to estimate $V^{*}_{\mathcal{R}}(s,u)$ for any RM task $\mathcal{R}$ over $\mathcal{AP}$, provided we can estimate $V^{*}_{\Diamond\varphi}(s)$ for any $\varphi$ in $\Diamond\mathbf{PL}(\mathcal{AP})$. However, the number of propositional formulas over $\mathcal{AP}$ (up to logical equivalence) is $2^{2^{|\mathcal{AP}|}}$ and modelling a separate OVF for each such task is intractable. To address this, we further decompose logical formulas based on their structure. Any formula $\varphi$ can be rewritten in disjunctive normal form, i.e., as a disjunction of conjunctions of literals (where a literal is either a proposition $x$ or its negation $\lnot x$). We then approximate the OVF of $\Diamond\varphi$ using the semantics of fuzzy logic [69], where $\max$ represents disjunction and $\min$ represents conjunction. [^2] While fuzzy operators have been widely applied in formal languages over quantitative signals [51, 52, 53, 54], we apply them here to approximate OVFs. To the best of our knowledge, this has not been previously explored.
Let $\varphi=\xi_{1}\lor\ldots\lor\xi_{k}$, where each $\xi_{i}$ is a conjunction of literals. We approximate:
$V^{*}_{\Diamond\varphi}(s)\approx\max_{i=1,\ldots,k}V^{*}_{\Diamond\xi_{i}}(s)$ (2)
For each conjunctive clause $\xi=l_{1}\land\ldots\land l_{k}$, where each $l_{i}$ is a literal, we approximate:
$V^{*}_{\Diamond\xi}(s)\approx\min_{i=1,\ldots,k}V^{*}_{\Diamond l_{i}}(s)$ (3)
By composing Approximations 13, we can estimate the OVF for any RM task based on only $2|\mathcal{AP}|$ OVFs—namely, those for $\Diamond x$ and $\Diamond\lnot x$, for each $x\in\mathcal{AP}$. We refer to these as the primitive value functions (PVFs).
6.3 Final Remarks
In this section, we showed that the optimal value function (OVF) of any RM task $\mathcal{R}$ can be approximated using just $2|\mathcal{AP}|$ primitive value functions (PVFs). From these $2|\mathcal{AP}|$ PVFs, we can estimate OVFs for doubly exponentially many logical tasks ($2^{2^{n}}$) and infinitely many RM tasks. Each PVF quantifies progress toward satisfying a single proposition or its negation, and can be learned directly from the grounding dataset $\mathcal{D}$ using any offline RL algorithm. This approach trades expressivity for modularity: directly modelling the OVFs of all RM tasks is infeasible, so we instead model a small, reusable set of $2|\mathcal{AP}|$ PVFs at the cost of introducing approximation error.
Leveraging that we can estimate the OVF for any RM task $\mathcal{R}$, we then extend Algorithm 1 via potential-based reward shaping. This allows us to provide a dense learning signal, even in environments that suffer from sparsity. Further details on learning PVFs, sources of approximation error, and the potential-based reward shaping scheme are provided in Appendix A.
Running Example.
Suppose our gridworld agent learns OVFs for the propositions $\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}},\mathrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}G}},\mathrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}},\triangle,\bigcirc$ and their negations. For propositions, these OVFs quantify progress toward reaching, for example, a red object or a circle. For negations, the OVFs capture progress toward reaching a state where the given proposition is false—e.g., any square without a red object. Note that these negations can typically be satisfied in a single step and their OVFs are trivial to model. These $2|\mathcal{AP}|$ OVFs are the primitive value functions (PVFs)—from these, the agent can estimate the OVF for any RM task.
7 Experiments
We conducted experiments to evaluate the following research questions:
  1. RQ1
    Grounded (Formal) Language Learning: Can formal language be grounded more effectively by leveraging compositional language structure?
  2. RQ2
    Compositional Generalization: Do grounded language interpretations generalize to out-of-distribution tasks beyond those observed in $\mathcal{D}$?
  3. RQ3
    Reinforcement Learning: Can agents use these grounded interpretations to autonomously learn new tasks through RL with self-generated rewards?
  4. RQ4
    Sparsity: Does our potential-based reward shaping strategy enable RL in sparse settings?
7.1 Experimental Setup
Figure 3: DrawerWorld (right) is a MuJoCo robotics environment where the agent can interact with two drawers and three boxes. Propositions capture whether each drawer is open, each box is lifted by the agent, and if a given box is in a given drawer.
We evaluated our approach on two domains: an image-based gridworld with randomized object locations called GeoGrid (introduced in our running example), and a continuous-control MuJoCo robotics environment called DrawerWorld (Figure 3). Full environment details are provided in Appendix B.1.
Grounding Datasets. To emulate realistic pretraining conditions, we collected the grounding dataset $\mathcal{D}$ under minimal assumptions about downstream tasks. In GeoGrid, we generated trajectories using a random policy. In DrawerWorld, we manually controlled the robot to perform generic behaviours (e.g., opening drawers, picking up boxes) under the constraint that each trajectory interacted with at most one box. We labelled these trajectories using a hand-crafted labelling function over a fixed vocabulary $\mathcal{AP}$. Note that such a labelling function is not a hard requirement—labelling could be done through other means such as human annotation or contrastive learning.
Tasks. We designed a diverse set of RM tasks (Table 1). To evaluate RQ2, we included tasks that involve behaviours rarely or never seen in $\mathcal{D}$: e.g., in DrawerWorld, the Pickup-Each-Box task requires interacting with all three boxes, whereas no single training trajectory involves interaction with more than one.
Table 1: List of RM tasks, with a description for each task. Tasks labelled (g) involve a temporally extended goal, tasks labelled (s) involve safety constraints, and tasks labelled (r) involve repeating some behaviour as often as possible. For each task, we report the mean and max (undiscounted) return across all trajectories in $\mathcal{D}$, along with the highest possible average return for any policy (or if unknown, we report the highest average return we are aware of, denoted with $\sim$). Some tasks involve behaviours that are rarely or never observed in $\mathcal{D}$.
Task Description Return in $\mathcal{D}$ Max Possible
Mean Max Return
GeoGrid
Sequence (g) Go to a red $\triangle$, then a green $\triangle$, then a blue $\triangle$. $0.04$ $1$ $1$
Loop (r) Repeatedly go to a red $\triangle$, then a green $\triangle$, then a blue $\triangle$. $0.04$ $3$ $\sim 5.36$
Logic (g+s) Go to all six objects, but always go to red objects before blue objects, and blue objects before green objects. $0.00$ $1$ $1$
Safety (g+s) Go to a red object, then a blue object, then a green object, but always avoid $\triangle$. $-0.84$ $1$ $1$
DrawerWorld
Lift-Red-Box (r) Lift and hold the red box as long as possible. $41.7$ $736$ $\sim 1538$
Pickup-Each-Box (g) Pick up the red box, then the blue box, then the green box. 00 00 $1$
Show-Green-Box (g) Reveal the green box if it’s in a closed drawer, then lift it. $0.22$ $1$ $1$
7.2 Method and Baselines
Ground-Compose-Reinforce (GCR). We implemented the full version of Ground-Compose-Reinforce with potential-based reward shaping. Both the labelling function $\hat{\mathcal{L}}$ and PVFs are neural networks trained from the fixed grounding dataset $\mathcal{D}$. In GeoGrid, we trained PVFs using Fitted Q-Iteration [70]. In DrawerWorld, we trained PVFs to predict Monte Carlo returns. [^3] We also evaluated MCQ [71], a state-of-the-art offline RL method, but it performed worse than Monte Carlo regression. We attribute this to the relatively small size of $\mathcal{D}$ compared to standard offline RL benchmarks.
Baselines. To answer RQ1-RQ3, we compared against the following end-to-end (i.e. non-compositional) baselines. Since these baselines do not use RMs to model temporal dependencies, they instead condition on past observations using a GRU [72]. See Appendix B.2 for complete details.
LTL-conditioned Behaviour Cloning (LTL-BC) is a neural network policy that directly maps LTL formulas to behaviours. We labelled each trajectory $\tau$ in $\mathcal{D}$ with a description $\varphi$ in LTL (based on the propositional labels for $\tau$ in $\mathcal{D}$), then trained the policy to maximize the log-likelihood of actions in $\tau$, conditioned on $\varphi$. For each downstream task in Table 1, we evaluated the policy by conditioning on an LTL formula that approximately aligns with that task. We also considered bespoke models trained with advance knowledge of the downstream tasks in Table 1. Bespoke Reward Model directly predicts rewards, optimal values, and terminations for all downstream tasks simultaneously, enabling RL with shaped rewards. To train this model, we labelled each trajectory $\tau$ in $\mathcal{D}$ with rewards and terminations for each downstream task based on the propositional labels in $\mathcal{D}$, then trained the model to directly predict these quantities. Optimal value estimates were trained via offline RL in a similar manner as GCR. Bespoke Behaviour Cloning (BC) is a neural network policy that directly imitates successful trajectories in $\mathcal{D}$ for every downstream task. Due to the limited number of reward-worthy trajectories in $\mathcal{D}$, we considered any trajectory achieving positive return on that task as successful.
To address RQ4, we compared various reward shaping schemes for GCR. No RS uses predicted RM rewards without any reward shaping (i.e. Algorithm 1). High-Level RS, inspired by Camacho et al. [46], uses a potential function that only considers the current RM state, but not the current MDP state.
Hyperparameters were tuned via a line search over batch size, learning rate, L1 regularization coefficient, and epochs (in that order), using a held-out 10% of the trajectories in $\mathcal{D}$. Final models were retrained on the full data. Full details of the experimental setup can be found in Appendix B.3.
7.3 Results
For each method (excluding behaviour cloning, which directly yields a policy), and for each task in Table 1, we trained a policy from scratch using PPO [73] (training details are reported in Appendix B.3). Recall that the agent does not observe ground-truth rewards or $\mathcal{L}^{*}$, and must generate its own learning signals. We report the performance of each policy at the end of training with respect to $\mathcal{L}^{*}$ (i.e. the ground-truth labelling function used to label $\mathcal{D}$) in Table 2. We measured performance as undiscounted return averaged over 100 episodes for GeoGrid or 20 episodes for DrawerWorld, over five independent RL runs. In Appendix B.4, we report RL learning curves, both with respect to the agent’s own reward model (without shaping rewards) and ground-truth rewards under $\mathcal{L}^{*}$.
GCR consistently solves novel RM tasks. Results show that GCR is the only method that consistently learns successful policies across both GeoGrid and DrawerWorld. It significantly outperforms even the best trajectories in $\mathcal{D}$ on Loop, Lift-Red-Box and Pickup-Each-Box (for the latter task, no successful demonstrations exist in $\mathcal{D}$). We attribute this to GCR’s ability to reuse its learned knowledge towards novel task compositions.
Non-compositional methods fail to generalize. Both LTL-BC, which is trained to imitate trajectories conditioned on entire LTL specifications, and bespoke methods, which directly predict task-specific rewards or behaviours given advance knowledge of the downstream tasks, fail to consistently produce reward-worthy behaviours. The learning curves show that Bespoke Reward Model assigns near-zero rewards to most trajectories in GeoGrid—likely due to the rarity of positive demonstrations in $\mathcal{D}$. In DrawerWorld, it produces highly misaligned rewards, resulting in a policy that achieves high return under its own reward model, but near-zero return under $\mathcal{L}^{*}$.
Reward shaping enables RL in sparse domains. Potential-based shaping provides modest improvements in GeoGrid, but is critical for success in DrawerWorld, where behaviours like opening a drawer and picking up a box are nearly impossible to discover with random exploration. The agent fails to learn under alternative reward shaping schemes (No RS, High-Level RS).
Table 2: Comparison of methods when used to train an RL policy on novel tasks using self-generated rewards. We report performance (undiscounted return, with respect to a ground-truth interpretation of each RM) at the end of training, averaged over 5 seeds with standard error.
Task GCR (Ours) LTL-BC Bespoke Bespoke No RS High-Level RS
Reward Model BC
GeoGrid
Sequence $\textbf{1.00}\pm 0.00$ $0.04\pm 0.01$ $0\pm 0$ $0.05\pm 0.01$ $0.94\pm 0.03$ $\textbf{1.00}\pm 0.00$
Loop $\textbf{5.36}\pm 0.08$ $0.03\pm 0.01$ $0\pm 0$ $0.04\pm 0.01$ $4.68\pm 0.05$ $5.27\pm 0.08$
Logic $\textbf{0.94}\pm 0.01$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $00.00\pm 0$ 0.00 $\textbf{0.94}\pm 0.01$
Safety $\textbf{1.00}\pm 0.00$ $-0.84\pm 0.01$ $-0.14\pm 0.01$ $-0.85\pm 0.01$ $0.23\pm 0.11$ $0.97\pm 0.01$
DrawerWorld
Lift-Red-Box $\textbf{1538}\pm 130$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$
Pickup-Each-Box $\textbf{1.00}\pm 0.00$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$
Show-Green-Box $\textbf{0.61}\pm 0.06$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$
8 Discussion and Limitations
In reflection of RQ1-RQ4, our experiments demonstrate the following advantages of our proposed Ground-Compose-Reinforce (GCR) approach. RQ1: GCR learns interpretations of formal language from data that are better aligned with the ground truth vs. end-to-end grounding approaches. RQ2: GCR generalizes to out-of-distribution tasks and behaviours that are rarely or never observed in the grounding dataset $\mathcal{D}$. RQ3: GCR allows an agent to learn new tasks via RL without supervision. RQ4: GCR with potential-based reward shaping effectively learns RL tasks in sparse settings. We now discuss the main limitations of this work, and opportunities for future work.
Extension to Other Languages. We study the GCR framework from the perspective of Reward Machines, however, we believe our core ideas apply to any formal language that admits compositional structure, such as those that deal with objects [74] and relations [75]. Extending our insights to systems interfaced via natural language (NL), which is often argued to be compositional [76], presents an important challenge. One promising approach is to autoformalize NL into formal language as an intermediate representation [34, 35, 36, 37]. In Appendix Section C, we show how modern LLMs like ChatGPT-4o and o3 (which do not need to be grounded in our specific environments) can zero-shot autoformalize NL reward function descriptions into RMs.
Extension to Other Problem Settings. Grounded language acquisition is a prerequisite for a myriad of language-conditioned problem settings. While we explore how grounded language can be leveraged for tasking RL agents without external reward supervision, future work could extend our insights to zero-shot execution of language tasks [5, 38, 44], question answering [77, 78], and interactive task learning [79, 80].
Reward Hacking. Misalignment between the agent’s interpretation of language and human intent can lead to unintended or harmful behaviours, posing potential societal risks if deployed. We aim to mitigate this through the use of a formal language, which is unambiguous over its propositional vocabulary. Still, ambiguity in the propositions themselves remains a concern. Future work could explore its impact, though prior works suggest the structure of formal languages may support more transparent reasoning under such uncertainty [55, 56].
Assumptions on Grounding Dataset $\mathcal{D}$. We assume the grounding dataset $\mathcal{D}$ is labelled with values of propositions. Future work could explore practical and scalable labelling methods, such as crowdsourced annotations [81] or self-supervised learning [82]. Our motivation also stems from the growing need for data-efficient methods of building AI systems with general capabilities. We show that even with limited data, our agent generalizes to diverse out-of-distribution tasks, a step toward building systems that require fewer resources, and lower environmental cost.
9 Conclusion
We present Ground-Compose-Reinforce, a framework for tasking RL agents directly via formal language. We propose a bottom-up approach to grounding language—first learning the meanings of individual symbols from data, then using formal compositional semantics to interpret complex tasks. We implement Ground-Compose-Reinforce for Reward Machines and show how any Reward Machine task can be grounded in dense, shaped rewards from a small grounding dataset of trajectories. Our experiments on challenging, sparse-reward tasks demonstrate its superior ability to interpret and solve novel tasks compared to grounding language directly from end to end.
Acknowledgements
We thank Harris Chan for his insightful and valuable input throughout all stages of this project. We gratefully acknowledge funding from the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Canada CIFAR AI Chairs Program. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute for Artificial Intelligence (https://vectorinstitute.ai/partnerships/). Finally, we thank the Schwartz Reisman Institute for Technology and Society for providing a rich multi-disciplinary research environment.
Supplementary Material for Ground-Compose-Reinforce: Tasking Reinforcement Learning Agents through Formal Language
In this supplementary material,
  • we further describe how we approximate optimal value functions and use that for potential-based reward shaping (Appendix A),
  • we provide experimental details (Appendix B),
  • we show how natural language descriptions of reward functions can be autoformalized into Reward Machines with modern LLMs like ChatGPT-4o (Appendix C),
  • and we discuss the potential societal impact of this work (Appendix D).
Appendix A Composition: Optimal Value Functions (Supplementary Details)
In this section, we
  • elaborate on Section 6.1’s description of how we approximate the optimal value function $V^{*}_{\mathcal{R}}(s,u)$,
    • first under the assumption that the RM does not contain self-loop transitions with non-zero rewards (Appendix A.1),
    • and then relaxing that assumption (Appendix A.2);
  • discuss approximation errors that can occur (Appendix A.3);
  • describe how we train with offline RL the primitive value functions used in the approximation (Appendix A.4);
  • and finally, show how we use the approximate value functions for potential-based reward shaping (Appendix A.5).
A.1 Approximating $V^{*}_{\mathcal{R}}(s,u)$ (No Rewarding RM Self-Transitions)
Algorithm 2 Value Iteration over RM States (Modified from Camacho et al. [46])
0:  RM $\mathcal{R}=\langle\mathcal{U},u_{0},\mathcal{F},\mathcal{AP},\delta_{u},\delta_{r}\rangle$, discount factor $\gamma_{\mathrm{RM}}$
1:  Initialize $v^{*}_{\mathcal{R}}(u)\leftarrow 0$, $\forall u\in\mathcal{U}$
2:  while not converged do
3:     for $u$ in $\mathcal{U}$ do
4:        $v^{*}_{\mathcal{R}}(u)\leftarrow\max_{\langle u^{\prime},\varphi,r\rangle\in\delta(u)}{\gamma_{\mathrm{RM}}(r+v^{*}_{\mathcal{R}}(u^{\prime}))}$
5:     end for
6:  end while
7:  return  $v^{*}_{\mathcal{R}}$
Recall that to estimate $V^{*}_{\mathcal{R}}(s,u)$, we evaluate each outgoing transition $\langle u,u^{\prime},\varphi,r\rangle$ from $u$ by combining the subtask value $V^{*}_{\Diamond\varphi}(s)$ with $v^{*}_{\mathcal{R}}(u^{\prime})$. While Camacho et al. [46] treat RM transitions as singular actions, we treat RM transitions as temporally extended options [83] that take a variable number of steps to execute. Specifically, for a transition $\langle u,u^{\prime},\varphi,r\rangle$, we make the following assumptions about the corresponding option:
  1. (1)
    It can be initiated if and only if the current RM state is $u$.
  2. (2)
    It optimizes for the subtask $\Diamond\varphi$ (i.e. satisfying $\varphi$ quickly and with high probability).
  3. (3)
    It either terminates when $\varphi$ is satisfied after a variable length of time $K$, or it never terminates.
  4. (4)
    The option will never result in a different transition in the RM than the one in question.
Our estimation makes use of the state-independent value function $v^{*}_{\mathcal{R}}(u)$, which we calculate using a variant of Value Iteration. Algorithm 2 shows the Value Iteration algorithm we use, modified from Camacho et al. [46] to reflect that rewards are garnered after the option completes. Note that Algorithm 2 should be run with a high-level discount factor $\gamma_{\mathrm{RM}}\ll\gamma$. This is because Algorithm 2 treats RM transitions as singular actions while in the actual RM-MDP, an RM transition may require many low-level steps to achieve. Under these conditions, the expected return of initiating the option corresponding to $\langle u,u^{\prime},\varphi,r\rangle$ in MDP state $s$ and RM state $u$ is:
$\mathbb{E}_{k\sim K,s^{\prime}}\bigl{[}\gamma^{k}(r+\gamma V^{*}_{\mathcal{R}}(s^{\prime},u^{\prime}))\bigr{]}$ $=\mathbb{E}_{s^{\prime}}\bigl{[}V^{*}_{\Diamond\varphi}(s)(r+\gamma V^{*}_{\mathcal{R}}(s^{\prime},u^{\prime}))\bigr{]}$
$\approx V^{*}_{\Diamond\varphi}(s)(r+\gamma v^{*}_{\mathcal{R}}(u^{\prime}))$
where $k\sim K$ represents the variable number of steps it takes for $\varphi$ to be satisfied (which might be $\infty$), and $s^{\prime}$ is the MDP state when the option terminates. We then estimate the optimal value function for RM task $\mathcal{R}$ by maximizing over the choice of option:
${V}^{*}_{\mathcal{R}}(s,u)$ $\approx\max_{\langle u^{\prime},\varphi,r\rangle\in\delta(u)}{\bigl{[}V^{*}_{\Diamond\varphi}(s)(r+\gamma v^{*}_{\mathcal{R}}(u^{\prime}))\bigr{]}}$
A.2 Approximating $V^{*}_{\mathcal{R}}(s,u)$ with Rewarding Self-Loop Transitions
To handle RMs that contain self-loop transitions with non-zero rewards, we require a few modifications. Intuitively, an option corresponding to a transition $\langle u,u^{\prime},\varphi,r\rangle$ where $u\neq u^{\prime}$ might receive a reward $r^{\prime}$ at any timestep it is active if there exists another transition $\langle u,u,\varphi^{\prime},r^{\prime}\rangle$. As a simple example, consider the RM in Figure 4. The RM describes a task where the agent needs to exit the lava as quickly as possible, and receives a reward of $-1$ for each step until it does so. When estimating the expected return of the edge labelled $\langle\lnot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}lava},0\rangle$, we need to consider the accumulation of the $-1$ reward at each timestep.
Figure 4: An RM that produces a reward of $-1$ for each timestep the agent spends in lava, until it exits the lava.
This can be hard to handle in general without more information, particularly when multiple self-loop edges exist for the same RM state. For simplicity, we assume that only non-self-loop transitions are treated as options (for the purposes of estimating optimal values), while all self-loop transitions in the current state $u$ are available at each step while an option is being executed. Thus, if we define $r_{u,u}$ as the maximum reward for any self-loop transition in RM state $u$, we can modify our expression for the expected return of an option corresponding to transition $\langle u,u^{\prime},\varphi,r\rangle$ in state $\langle s,u\rangle$ as follows:
$\mathbb{E}_{k\sim K,s^{\prime}}\bigl{[}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r_{u,u}+\gamma r_{u,u}+\ldots+\gamma^{k-1}r_{u,u}}+\gamma^{k}(r+\gamma V^{*}_{\mathcal{R}}(s^{\prime},u^{\prime}))\bigr{]}$
$=\mathbb{E}_{k\sim K,s^{\prime}}\Bigl{[}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r_{u,u}\Bigl{(}\frac{1-\gamma^{k}}{1-\gamma}\Bigr{)}}+\gamma^{k}(r+\gamma V^{*}_{\mathcal{R}}(s^{\prime},u^{\prime}))\Bigr{]}$
$=\mathbb{E}_{s^{\prime}}\Bigl{[}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r_{u,u}\Bigl{(}\frac{1-V^{*}_{\Diamond\varphi}(s)}{1-\gamma}\Bigr{)}}+V^{*}_{\Diamond\varphi}(s)(r+\gamma V^{*}_{\mathcal{R}}(s^{\prime},u^{\prime}))\Bigr{]}$
$\approx{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r_{u,u}\Bigl{(}\frac{1-V^{*}_{\Diamond\varphi}(s)}{1-\gamma}\Bigr{)}}+V^{*}_{\Diamond\varphi}(s)(r+\gamma v^{*}_{\mathcal{R}}(u^{\prime}))$
For the previous lava example, the expected return for the option that corresponds to the transition labelled $\langle\lnot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}lava},0\rangle$ now correctly reflects a reward of $-1$ obtained for each step until the option terminates by the agent exiting the lava. We modify our approximation of ${V}^{*}_{\mathcal{R}}(s,u)$ as follows:
${V}^{*}_{\mathcal{R}}(s,u)$ $\approx\max_{\langle u^{\prime},\varphi,r\rangle\in\delta(u)}{\Bigl{[}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r_{u,u}\Bigl{(}\frac{1-V^{*}_{\Diamond\varphi}(s)}{1-\gamma}\Bigr{)}}+V^{*}_{\Diamond\varphi}(s)(r+\gamma v^{*}_{\mathcal{R}}(u^{\prime}))\Bigr{]}}$ (4)
We also modify the Value Iteration algorithm to consider self-loops with non-zero rewards (see Algorithm 3).
Algorithm 3 Value Iteration over RM States (modified for self-loop transitions)
0:  RM $\mathcal{R}=\langle\mathcal{U},u_{0},\mathcal{F},\mathcal{AP},\delta_{u},\delta_{r}\rangle$, MDP discount factor $\gamma$, high-level discount factor $\gamma_{\mathrm{RM}}$
1:  Initialize $v^{*}_{\mathcal{R}}(u)\leftarrow 0$, $\forall u\in\mathcal{U}$
2:  Extract maximum self-loop rewards $r_{u,u}$, $\forall u\in\mathcal{U}$ from $\delta_{r}$
3:  while not converged do
4:     for $u$ in $\mathcal{U}$ do
5:        $v^{*}_{\mathcal{R}}(u)\leftarrow\max_{\langle u^{\prime},\varphi,r\rangle\in\delta(u)}{\bigl{(}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}r_{u,u}\bigl{(}\frac{1-\gamma_{\mathrm{RM}}}{\gamma}\bigr{)}}+\gamma_{\mathrm{RM}}(r+v^{*}_{\mathcal{R}}(u^{\prime}))\bigr{)}}$
6:     end for
7:  end while
8:  return  $v^{*}_{\mathcal{R}}$
A.3 Sources of Approximation Errors
We now discuss how errors can occur when estimating $V^{*}_{\mathcal{R}}(s,u)$ from PVFs.
First, Approximation 4 clearly introduces approximation error from assuming the largest reward among self-loop transitions $r_{u,u}$ will be garnered until an outgoing transition is reached, and from using a bootstrapped value estimate of the next state $v^{*}_{\mathcal{R}}(u^{\prime})\approx V^{*}_{\mathcal{R}}(s^{\prime},u^{\prime})$. Another less obvious source of error is that the expected return for each outgoing transition $\langle u,u^{\prime},\varphi,r\rangle$ is estimated by assuming that an optimal policy for reaching $\varphi$ will be followed. However, this ignores the possibility of some other transition from $u$ occurring before the intended transition. In general, it may not be possible to satisfy $\varphi$ while achieving value $V^{*}_{\Diamond\varphi}(s)$ in the subtask $\Diamond\varphi$ while also avoiding all other transitions from RM state $u$ that are not $\langle u,u^{\prime},\varphi,r\rangle$.
Estimating the OVF for a disjunction of formulas via Approximation 2 always underestimates the true value. We prove this as follows. Suppose $\varphi=\xi_{1}\lor\ldots\lor\xi_{k}$ and recall that the approximation of $V^{*}_{\Diamond\varphi}(s)$ is $\max_{i=1,\ldots,k}V^{*}_{\Diamond\xi_{i}}(s)$. Observe that for each $i$, $V^{*}_{\Diamond\xi_{i}}(s)\leq V^{*}_{\Diamond\varphi}(s)$, since for every trajectory $\tau$, if $\xi_{i}$ is satisfied at timestep $T$ in $\tau$, then $\varphi$ must also be satisfied at timestep $T$ (or earlier) in $\tau$. Thus, $\max_{i=1,\ldots,k}V^{*}_{\Diamond\xi_{i}}(s)\leq V^{*}_{\Diamond\varphi}(s)$. The reason this bound is not tight is because there are situations where satisfying one of $\xi_{1},\ldots,\xi_{k}$ is easier than satisfying any of $\xi_{1},\ldots,\xi_{k}$ individually. For example, consider the task $\Diamond(X\lor\lnot X)$, which is always trivially solved on the first step. However, it is possible that $\Diamond X$ and $\Diamond\lnot X$ are both non-trivial tasks.
Lastly, estimating the OVF for a conjunction of formulas via Approximation 3 always overestimates the true value, by a similar line of reasoning as for disjunction. In general, knowing $V^{*}_{\Diamond\xi_{1}}(s)$ and $V^{*}_{\Diamond\xi_{2}}(s)$ does not provide enough information to estimate $V^{*}_{\Diamond(\xi_{1}\land\xi_{2})}(s)$. For instance, it may be the case that $\xi_{1},\xi_{2}$ are mutually exclusive and thus, $V^{*}_{\Diamond(\xi_{1}\land\xi_{2})}$ is zero everywhere. However, whether or not $\xi_{1},\xi_{2}$ are mutually exclusive cannot be inferred based only on $V^{*}_{\Diamond\xi_{1}}(s)$ and $V^{*}_{\Diamond\xi_{2}}(s)$. Nangue Tasse et al. [84] provide a discussion on this topic for a related setting.
A.4 Training Primitive Value Functions
PVFs can be trained directly from the grounding dataset $\mathcal{D}$ based on any offline RL approach. Algorithm 4 shows how to train the PVF for a single primitive task $\Diamond x$ using a simple offline RL algorithm (Fitted Q-Iteration [70]). The approach can be easily adapted to negations $\Diamond\lnot x$ as well, and in practice, we simultaneously train all $2|\mathcal{AP}|$ possible PVFs in parallel as a single neural network.
Algorithm 4 Learning PVF $V^{*}_{\Diamond_{x}}(s)$ for $x\in\mathcal{AP}$ from Trajectory Data $\mathcal{D}$
0:  Dataset $\mathcal{D}=\left\{\langle\tau^{i},\omega^{i}\rangle\right\}_{i=1}^{N}$, discount factor $\gamma$, proposition $x\in\mathcal{AP}$
1:  Initialize Q function $Q_{\theta}:\mathcal{S}\times\mathcal{A}\to\mathbbm{R}$
2:  Initialize value function $V_{\psi}:\mathcal{S}\to\mathbbm{R}$
3:  while not converged do
4:     Sample transition $\langle s,\omega,a,s^{\prime},\omega^{\prime}\rangle\sim\mathcal{D}$
5:     Update $\phi$ with SGD on $\mathrm{BinaryCrossEntropy}(\mathcal{L}_{\phi}(s),\mathbbm{1}[x\in\omega])$
6:     $\mathrm{reward}\leftarrow\mathbbm{1}\left[x\in\omega^{\prime}\right],\mathrm{next\_value}\leftarrow\max_{a^{\prime}\in\mathcal{A}}Q_{\theta}(s^{\prime},a^{\prime}),\mathrm{done}\leftarrow\mathbbm{1}\left[x\in\omega^{\prime}\right]$
7:     Update $\theta$ with SGD on $\bigl{(}Q_{\theta}(s,a)-\mathrm{stop\_grad}(\mathrm{reward}+\gamma*(1-\mathrm{done})*\mathrm{next\_value})\bigr{)}^{2}$
8:     Update $\psi$ with SGD on $(V_{\psi}(s^{\prime})-\mathrm{next\_value})^{2}$
9:  end while
10:  return  $V_{\psi}$
A.5 Potential-based Reward Shaping
We extend our core Ground-Compose-Reinforce algorithm with potential-based reward shaping by leveraging Approximations 1-3 and trained PVFs to predict $V^{*}_{\mathcal{R}}(s,u)$ for any MDP state $s$ and RM state $u$. This is shown in Algorithm 5 with changes from the core algorithm highlighted in red.
Algorithm 5 Ground-Compose-Reinforce for RMs with Potential-Based Reward Shaping
0:  MDP $\mathcal{M}$ without rewards, Propositional symbols $\mathcal{AP}$, Dataset $\mathcal{D}$ of labelled trajectories, RM task $\mathcal{R}$ over $\mathcal{AP}$, Shaping potential weighting coefficient $\lambda$ // Phase 1: Ground symbols via offline learning
1:  Train labelling function ^$\hat{\mathcal{L}}(s)$ on $\mathcal{D}$ using any binary classifier
2:  Train PVFs $V^{*}_{\Diamond x}(s)$ and $V^{*}_{\Diamond\lnot x}(s)$, $\forall x\in\mathcal{AP}$ on $\mathcal{D}$ // Phase 2: Solve any RM task using self-generated rewards
3:  Initialize policy $\pi_{\mathcal{R}}(a\mid s,u)$ arbitrarily
4:  for each episode do
5:     Observe initial state $s$ in $\mathcal{M}$; set $u$ to the initial state of $\mathcal{R}$
6:     Estimate initial value $v\approx V^{*}_{\mathcal{R}}(s,u)$ using Approximations 1-3 and trained PVFs
7:     while $u$ is non-terminal do
8:        Sample action $a\sim\pi_{\mathcal{R}}(\cdot\mid s,u)$
9:        Execute $a$ in $\mathcal{M}$ and observe next state $s^{\prime}$
10:        Compute truth assignment $\hat{\omega}\leftarrow\hat{\mathcal{L}}(s^{\prime})$
11:        Update RM: $u^{\prime}\leftarrow\delta_{u}(u,\hat{\omega})$, $r\leftarrow\delta_{r}(u,\hat{\omega})$
12:        Update value $v^{\prime}\approx V^{*}_{\mathcal{R}}(s^{\prime},u^{\prime})$ using Approximations 1-3 and trained PVFs
13:        Update policy $\pi_{\mathcal{R}}$ with RL for transition $\langle s,u,a,r+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\lambda(\gamma v^{\prime}-v)},s^{\prime},u^{\prime}\rangle$
14:        Set $s\leftarrow s^{\prime}$, $u\leftarrow u^{\prime},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v\leftarrow v^{\prime}}$
15:     end while
16:  end for
Appendix B Experimental Details
B.1 Domain Descriptions
Environments.
We considered two novel environments. GeoGrid is an $8\times 8$ image-based gridworld depicted in Figure 1 with six objects randomly positioned at the start of each episode. States are $8\times 8\times 6$-dimensional images that identify each cell’s colour/shape and the agent’s location, while propositions $\{\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}},\mathrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}G}},\mathrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}},\triangle,\Circle\}$ identify if the agent is at an object with that particular colour or shape. DrawerWorld is a MuJoCo environment adapted from MetaWorld [85]. The agent controls a robotic gripper and can interact with two drawers (left and right) and three boxes (red, green, and blue). Observations are 78-dimensional vectors representing positions of objects and the gripper. Propositions identify whether a particular drawer is open, whether a particular block is picked up by the agent, and whether a particular block is currently inside a particular drawer.
Datasets.
We carefully curated datasets $\mathcal{D}$ in each environment to support our analysis of compositional generalization (RQ2). In GeoGrid, $\mathcal{D}$ is comprised of 5000 trajectories of length 100 generated under a random-action policy. In DrawerWorld, $\mathcal{D}$ is comprised of 350 trajectories of varying length that we collected by manually controlling the robot in the simulator. To ensure sufficient state coverage in $\mathcal{D}$, drawers and boxes were initialized in a random configuration when collecting each trajectory. Behaviours that appear in the dataset include opening and closing drawers, picking up boxes, and moving boxes from one location to another, but no trajectory involves direct interaction with more than one box. A small number of trajectories involve incidental (but not prolonged) interaction with more than one box (e.g. bumping into one box while moving another).
Tasks.
We designed a diverse set of RM tasks (Table 1). For tasks that involve achieving a (temporally extended) goal, the RM terminates and provides a reward of 1 upon doing so. The RM terminates with a reward of 0 if the goal becomes logically impossible (e.g. due to breaking a constraint), except in Safety, which terminates with a penalty of $-1$. Loop and Lift-Red-Box involve repeating some desired behaviour, and the RM yields a reward of 1 for each such repetition.
B.2 Baseline Implementation Details
All approaches involve supervised training on $\mathcal{D}$, and Ground-Compose-Reinforce and Bespoke Reward Model additionally require an RL phase in the environment. Network architectures for supervised training on $\mathcal{D}$ are reported in Table 7. PPO network architectures are reported in Table 8. All policy networks (whether trained via RL or behaviour cloning) use GRUs to temporal dependencies except for GCR, which uses RM transitions. The policy network outputs a probability distribution over actions (in DrawerWorld, the outputs of the network parameterize a Gaussian policy’s mean and standard deviation). For methods relying on potential-based reward shaping, we computed shaping rewards without discounting the next potential, i.e. we issued the shaping reward as $\lambda(v^{\prime}-v)$ rather than $\lambda(\gamma v^{\prime}-v)$. Though this loses some theoretical convergence properties, we found it to significantly outperform the standard shaping reward in all cases.
Ground-Compose-Reinforce. GCR consists of the following neural networks: a labelling function network that outputs a single binary classification logit for each proposition in $\mathcal{AP}$, a PVF network that outputs an optimal value prediction for each literal in $\mathcal{AP}$, and a policy of the form $\pi(a_{t}|s_{t},u_{t})$ that conditions on the current RM state. The labelling function is trained via a binary cross entropy loss on $\mathcal{D}$, the PVFs are trained via offline RL on $\mathcal{D}$, and the policy is trained via RL supported by the labelling function and PVFs to provide learning signals.
LTL-conditioned Behaviour Cloning. This baseline models a neural network policy $\pi_{\theta}(a_{t}|h_{t},\varphi)$, where the history $h_{t}$ is encoded by a GRU [72] and $\varphi$ (a goal represented directly in LTL) is encoded by a Transformer [86]. Only observations (and not actions) are encoded as part of the history. We trained $\pi_{\theta}$ by labelling each trajectory $\tau^{i}$ in $\mathcal{D}$ with an LTL formula $\varphi^{i}$ based on the sequence of propositional labels $\sigma^{i}$ in $\mathcal{D}$ and then minimized the behaviour cloning loss $\mathbb{E}_{\mathcal{D}}[-\log{\pi_{\theta}(a_{t}|h_{t},\varphi)}]{}$. Finally, we evaluated the model on the downstream tasks by conditioning on the LTL formulas shown in Table 3.
To the best of our knowledge, there are no existing approaches that generate LTL descriptions based on a single trajectory. We instead used a custom approach based on common specification templates to generate diverse LTL descriptions. For each trajectory $\tau$ in $\mathcal{D}$, we randomly generated a single formula that is satisfied by $\tau$ for each of the following specification templates found in Table 2 of Menghi et al. [87]: visit, sequenced visit, ordered visit, patrolling (for this purpose, we consider an event to occur infinitely often if it occurs at least five times within the same trajectory in GeoGrid or 200 times within the same trajectory in DrawerWorld) and global avoidance. These templates were chosen since they correspond to LTL properties that are relatively simple to automatically mine from a given trajectory. We then labelled each trajectory $\tau$ with a randomly chosen LTL formula from among this set.
Bespoke Reward Model. This baseline is a single neural network that directly predicts the reward, termination, and optimal value function for each of the downstream tasks. The neural network consists of an observation encoder, followed by two GRU layers (to encode the history of observations), followed by three linear output heads to predict rewards, optimal values, and terminations, respectively, for all downstream tasks simultaneously for that domain. To generate target rewards and terminations for a trajectory $\tau$ in $\mathcal{D}$, we evaluated the RM of each downstream task based on the sequence of propositional labels $\sigma^{i}$. The optimal value estimates were trained using offline RL in a similar manner as the PVFs. Finally, a policy was obtained using RL on the rewards and terminations, while the optimal values were used for potential-based reward shaping, similar to the shaped version of GCR.
Bespoke Behaviour Cloning. This baseline is similar to LTL-BC, except it does not condition on an LTL task—instead, it simultaneously outputs actions for each of the possible downstream tasks. To evaluate the policy on a specific downstream task, only the output for that task is considered. We trained the policy via behaviour cloning on any trajectory that achieves positive return on a particular downstream task due to the limited number of successful demonstrations.
Table 3: LTL formulas used to evaluate LTL-BC.
Task LTL Formula
GeoGrid (propositions are named r,g,b,c,t instead of $\mathrm{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}R}},\mathrm{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}G}},\mathrm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}},\triangle,\Circle$ to avoid confusion with LTL operators)
Sequence $\Diamond\bigl{(}(r\wedge t)\wedge\Diamond((g\wedge t)\wedge\Diamond(b\wedge t))\bigr{)}$
Loop $\Box\Diamond\bigl{(}(r\wedge t)\wedge\Diamond((g\wedge t)\wedge\Diamond(b\wedge t))\bigr{)}$
Logic $\Diamond(r\wedge t)\wedge\Diamond(g\wedge t)\wedge\Diamond(b\wedge t)\wedge\Diamond(r\wedge c)\wedge\Diamond(g\wedge c)\wedge\Diamond(b\wedge c)\wedge(\neg b\,\mathcal{U}\,r)\wedge(\neg g\,\mathcal{U}\,b)$
Safety $\Diamond\bigl{(}(r\wedge t)\wedge\Diamond((g\wedge t)\wedge\Diamond(b\wedge t))\bigr{)}\wedge\Box\neg t$
DrawerWorld
Lift-Red-Box $\Box\Diamond\mathrm{RedBoxLifted}$
Pickup-Each-Box $\Diamond(\mathrm{RedBoxLifted}\wedge\Diamond(\mathrm{BlueBoxLifted}\wedge\Diamond\mathrm{GreenBoxLifted}))$
[…]
B.3 Experimental Setup and Hyperparameter Details
Details: Supervised Training on $\mathcal{D}$.
All experiments were run on a compute cluster. Supervised training on $\mathcal{D}$ required a single GPU and CPU, minimal memory resources (24GB of RAM or less) and no more than 30 minutes to train any method to 100 epochs. We tuned hyperparameters via a line search over batch size, learning rate, L1 regularization coefficient, and epochs (in that order) using a held-out 10% of the trajectories in D, and the final hyperparameters are reported in Table 5. Final models were retrained on the full data with the tuned hyperparameteres. We note that the batch size hyperparameter should be interpreted differently for methods requiring a GRU. For GCR, it refers to the number of transitions sampled from $\mathcal{D}$. For LTL-BC, Bespoke Reward Model and Bespoke Behaviour Cloning, it refers to the number of full length trajectories sampled from $\mathcal{D}$. This is because it is necessary to keep transitions in a trajectory in the correct order to train the GRU.
Details: RL Training.
All experiments were run on a compute cluster. Each RL run used a single GPU, 16 CPUs, and 48GB of RAM. For GCR, runs took up to 6 hours on GeoGrid (to train to 15M frames) and 16 hours on DrawerWorld (to train to 20M frames). RL training with the Bespoke Reward Model took longer due to GRUs—up to 12 hours on GeoGrid (to train to 15M frames) and 18 hours on DrawerWorld (to train to 20M frames). For RL training, we used the implementation of PPO at https://github.com/lcswillems/torch-ac with the hyperparameters in Table 6. The total number of environment steps each method was trained on was 2.5M for Sequence, 4M for Loop, 10M for Safety, 20M for Show-Green-Box, and 15M for all others.
B.4 Learning Curves
We report RL learning curves for each method and task in Figures 5 and 6. GCR and its variants have an internal reward model that is better aligned with the ground truth compared to Bespoke Reward Model. In GeoGrid, GCR’s reward model is near perfect, and in all cases, optimizing its internal rewards improves ground truth performance as well. Bespoke Reward Model almost always predicts near-zero rewards on GeoGrid tasks (except Safety), likely since there are few examples of positive demonstrations in $\mathcal{D}$. On DrawerWorld, Bespoke Reward-Model is highly misaligned, predicting large rewards but garnering near-zero return based on the ground truth.
In terms of sample efficiency, we observe that GCR outperforms all baselines. The difference is marginal in GeoGrid, where exploration is less of an issue, but in DrawerWorld, all other reward shaping approaches fail.
Appendix C Autoformalizing Natural Language to Reward Machines
While several works have been dedicated to the autoformalization of natural language instructions into formal languages such as LTL [34, 35, 36, 37], we show that a modern large language model can perform this task zero-shot for RMs, without having specifically been trained on it (to the best of our knowledge). This allows us to directly task the RL agent in our framework through natural language, autoformalize the task description into an RM, and train the agent using its grounded formal language understanding. Critically, the autoformalizer’s understanding of language does not need to be grounded.
We tested OpenAI’s ChatGPT-4o and o3 models as the autoformalizer and prompted it with a description of the autoformalization task (including the output format), a text description of the environment, a list of propositions and associated text descriptions, and a text description of the desired reward function (Listing 1). For each of the four GeoGrid tasks in Table 1, we ran each autoformalizer five times for consistency. We manually evaluated each outputted RM based on whether it yielded a reward function that exactly matched the textual description, with the success rate reported in Table 4.
ChatGPT-4o correctly produced RMs for all tasks, except for Logic, which requires a complex RM (our solution involved 10 states and 18 transitions). However, we note that ChatGPT-4o was nearly correct for all five trials for Logic—each of its outputted RMs deviated by a single transition that changed the behaviour of the resultant reward function. o3 outputted correct RMs on all tasks. Notably, the outputted RMs were identical in structure to the intended RMs we manually constructed in all cases (with the only differences being in the numbering of RM states and the representation of equivalent logical formulas).
You are given a list of propositional symbols and their descriptions, an environment description, and a description of a desired reward function in English. Your job is to construct a Reward Machine representing this reward function.
[…]
Environment Description: The environment is a gridworld, where some squares have objects. Each object has a colour (red, green, or blue) and a shape (circle, or triangle).
Propositions:
- red: The agents current cell has a red object.
- blue: The agents current cell has a blue object.
- green: The agents current cell has a green object.
- triangle: The agents current cell has a triangle.
- circle: The agents current cell has a circle.
Task: <TASK DESCRIPTION>
Listing 1: Reward Machine Autoformalization Prompt
Table 4: Success rate for ChatGPT-4o and o3 when autoformalizing Reward Machines from a natural language description of the desired reward function.
Task Description GPT-4o Success Rate o3 Success Rate
Sequence Give a reward of 1 and terminate the episode when a red triangle, a green triangle, and a blue triangle have been reached, in that order. Only give the reward of 1 when the final step has been completed. Give 0 reward and never terminate the episode otherwise. 100% 100%
Loop Give a reward of 1 when a red triangle, a green triangle, and a blue triangle have been reached, in that order. Only give the reward of 1 when the final step has been completed. After completing this sequence, the agent may repeat all steps of the sequence to receive the reward again, as many times as it wishes. The episode never terminates. 100% 100%
Logic Give a reward of 1 and terminate the episode as soon as a red triangle, red circle, green triangle, green circle, blue triangle, and blue circle have all been reached at some point. However, blue objects should not be visited until both red objects are visited, and green objects should not be visited until both blue objects are visited. Circles and triangles of the same colour can be reached in either order. If any objects are reached out of order, immediately terminate the episode with a reward of 0. 0% 100%
Safety Give a reward of 1 when a red object, a green object, and a blue object have been reached, in that order. Only give the reward of 1 when the final step has been completed. However, always avoid squares with triangles—if this is violated, immediately terminate the episode with a reward of -1. 100% 100%
Appendix D Societal Impact
Data-efficient learning lowers environmental cost. Ground-Compose-Reinforce (GCR) achieves strong generalization from a relatively small, task-agnostic trajectory dataset. Because it avoids the need for internet-scale demonstrations, the total compute and data collection burden is substantially reduced, which in turn diminishes the carbon footprint of training and retraining large decision-making systems.
Transparent, verifiable task specifications. By exposing an explicit formal specification layer (Reward Machines) between the human and the agent, GCR allows auditors to read, simulate, and formally verify the reward logic before deployment. This contrasts with opaque end-to-end reward models and can help regulators trace undesirable behaviour back to a concrete symbolic condition rather than a latent neural representation, supporting safer and more accountable RL pipelines.
Broader access to capable agents. Because symbols are grounded once and then recomposed to create an unbounded task space, domain experts without ML backgrounds can author complex tasks simply by writing formulas, potentially democratizing advanced robotics and simulation tools in education, manufacturing, and assistive settings. The same mechanism lets small-lab researchers prototype complex multi-stage tasks without the costs associated with collecting new labelled rewards.
Reward hacking and specification gaps. If the learned grounding model assigns the wrong propositions to some states, the formal specification no longer matches human intent. In the paper, we caution that such mis-grounding can lead to “unintended or harmful behaviours” despite the use of a precise formal language. Attackers could also deliberately craft adversarial inputs that flip propositional truths so that the agent believes it has satisfied (or avoided) certain sub-goals, creating safety hazards.
Misuse through automated physical systems. Because GCR enables non-experts to compose lengthy, conditional task logic, it could accelerate deployment of autonomous robots in manufacturing, warehousing, or surveillance. In the absence of stringent safety checks, these robots might execute hazardous motion plans or be repurposed for malign objectives (e.g., automating the handling of dangerous substances or aiding illicit surveillance).
Labour displacement. Easier programming of general-purpose robotic skills may substitute for manual labour in logistics or assembly lines, contributing to job displacement without adequate social safety nets.
Table 5: Hyperparameters for Supervised Training on $\mathcal{D}$.
Hyperparameter GCR LTL-BC Bespoke Reward Model Bespoke BC
Labelling Function PVFs
GeoGrid
Batch size 256 1024 100 100 100
Learning rate 3e-4 3e-4 1e-4 3e-3 3e-4
L1 loss 1e-5 0 0 0 0
Epochs 10 100 9 100 11
Discount factor n.a. 0.97 n.a. 0.97 n.a.
DrawerWorld
Batch size 256 256 50 50 50
Learning rate 3e-4 3e-4 1e-3 1e-4 1e-4
L1 loss 1e-5 1e-5 0 0 1e-9
Epochs 100 100 87 11 4
Discount factor n.a. 0.9975 n.a. 0.9975 n.a.
Table 6: RL Training Hyperparameters.
Hyperparameter Value for all methods
GeoGrid
Number of parallel environments 16
Frames per update per process 1000
Learning rate 3e-4
Discount factor ($\gamma$) 0.97
GAE parameter ($\lambda$) 0.95
Clip range 0.2
Entropy coefficient 1e-4
Value loss coefficient 0.5
Number of epochs per update 4
Minibatch size 4000
[…]
Shaping potential weighting coefficient ($\lambda$) 1
DrawerWorld
Number of parallel environments 16
Frames per update per process 4000
Learning rate 3e-4
Discount factor ($\gamma$) 0.99
GAE parameter ($\lambda$) 0.99
Clip range 0.2
Entropy coefficient 0.01 in Pickup-Each-Box, otherwise 0.03
Value loss coefficient 0.5
Number of epochs per update 10
Minibatch size 8000
[…]
Shaping potential weighting coefficient ($\lambda$) 0.1 in Lift-Red-Box, otherwise 1
Table 7: Network Architectures for Supervised Training on $\mathcal{D}$.
GCR LTL-BC Bespoke Reward Model Bespoke BC
Labelling Function PVFs
GeoGrid
Conv2d(6,16,3,1,1) ReLU Conv2d(16,32,3,1,1) ReLU Flatten Linear(2048,128) ReLU Linear(128,5) Conv2d(6,32,3,1,1) ReLU Conv2d(32,32,3,1,1) ReLU Flatten Linear(2048,256) ReLU Linear(256,10) Obs Encoder: Conv2d(6,16,3,1,1) ReLU Conv2d(16,32,3,1,1) ReLU Flatten Linear(2048,256) LTL Encoder: Transformer(d_model=64, nhead=4, dim_feedforward=128, num_layers=2) Policy: GRU(256+64,256) ReLU GRU(256,256) Linear(768,16) Conv2d(6,16,3,1,1) ReLU Conv2d(16,32,3,1,1) ReLU Flatten Linear(2048,256) GRU(256,256) ReLU GRU(256,256) Linear(768,4) $\times$ 3 Conv2d(6,16,3,1,1) ReLU Conv2d(16,32,3,1,1) ReLU Flatten Linear(2048,256) GRU(256,256) ReLU GRU(256,256) Linear(768,16)
DrawerWorld
Linear(39,1600) ReLU Linear(1600,11) Linear(39,1600) ReLU Linear(1600,400) ReLU Linear(400,22) Obs Encoder: Linear(39,1600) ReLU Linear(1600,400) LTL Encoder: Transformer(d_model=64, nhead=4, dim_feedforward=128, num_layers=2) Policy: GRU(464,256) ReLU GRU(256,256) Linear(976,4) Linear(39,1600) ReLU Linear(1600,400) ReLU GRU(400,256) ReLU GRU(256,256) Linear(912,3) $\times$ 3 Linear(39,1600) ReLU Linear(1600,400) ReLU GRU(400,256) ReLU GRU(256,256) Linear(912,24)
Table 8: Network Architectures for PPO.
GCR Bespoke Reward Model
GeoGrid
Encoder: Conv2d(6,16,3,1,1) ReLU Conv2d(16,32,3,1,1) ReLU Flatten Actor Head: Linear(2048 + $|\mathcal{U}|$,128) ReLU Linear(128,64) ReLU Linear(64,4) Critic Head: Linear(2048 + $|\mathcal{U}|$,128) ReLU Linear(128,64) ReLU Linear(64,1) Encoder: Conv2d(6,16,3,1,1) ReLU Conv2d(16,32,3,1,1) ReLU Flatten GRU(2048,128) Actor Head: Linear(128,128) ReLU Linear(128,64) ReLU Linear(64,4) Critic Head: Linear(128,128) ReLU Linear(128,64) ReLU Linear(64,1)
DrawerWorld
Actor: Linear(78 + $|\mathcal{U}|$,512) ReLU Linear(512,512) ReLU Linear(512,512) ReLU Linear(512,8) Critic: Linear(78 + $|\mathcal{U}|$,512) ReLU Linear(512,512) ReLU Linear(512,512) ReLU Linear(512,1) Encoder: Linear(78,512) ReLU Linear(512,512) GRU(512,512) Actor Head: Linear(512,512) ReLU Linear(512,8) Critic Head: Linear(512,512) ReLU Linear(512,1)
Figure 5: RL learning curves for GeoGrid tasks, showing returns under the agent’s own reward model (“perceived rewards”) as well as rewards under the ground-truth $\mathcal{L}^{*}$ (“actual rewards”). Error bars show standard error.
Figure 6: RL learning curves for DrawerWorld tasks, showing returns under the agent’s own reward model (“perceived rewards”) as well as rewards under the ground-truth $\mathcal{L}^{*}$ (“actual rewards”). Error bars show standard error.